OpenStudy (anonymous):

f(x)=5cos(7ln(x)) , find f(x)`

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

you need its derivative ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\Large f(x)=5\cos(7\ln(x))\] using standard formulas with chain rule . \[\Large f'(x)=-5\sin(7(\ln(x))*\frac{d}{dx}(7\ln(x)\] \[\Large -35\sin(7\ln(x))*\frac{1}{x}\]

OpenStudy (anonymous):

i entered it , and it ws wrong

OpenStudy (anonymous):

(online homework)

jimthompson5910 (jim_thompson5910):

post a screenshot of what you entered

OpenStudy (anonymous):

ok.

jimthompson5910 (jim_thompson5910):

you're missing a 35

jimthompson5910 (jim_thompson5910):

otherwise, everything else is correct

jimthompson5910 (jim_thompson5910):

you'll probably want a decimal approximation for the second one

OpenStudy (anonymous):

Ops i did miss out 35 . its correct! thanks I wont to give both of medal. how do i do that?

jimthompson5910 (jim_thompson5910):

just give it to sami since that user did all the work really